A function f is given by f(4x-1) = 4x2+5. Evaluate f(15).
Answer by Theo(13342)

i don't think i ever saw oone like this, but i'll give it a shot.

your problem is f(4x - 1) = 4x^2 + 5

you want to know what f(15) is?

if it were f(x) = 4x^2 + 5, you would replace x = 15 and get f(15) = 4 * 15^2 + 5 which would become f(15) = 4 * 225 + 5 which would become f(15) = 905.

since it's f(4x-1), i believe you would set 4x-1 = 15 and solve for x to get x = 4.

f(4x-1) = 4x^2 + 5 would become f(15) = 4 * 4^2 + 5 which would become f(15) = 4 * 16 + 5 which would become f(15) = 64 + 5 which would become f(15) = 69.

this seems logical to me and i did see some references on the web that would indicate the logic is good.